OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 5(x + 2) greater than or equal to 7x + 2(x - 1)? A number line with a closed circle on 3 and shading to the right A number line with a closed circle on 3 and shading to the left 3 greater than or equal to x x less than or greater to 3

OpenStudy (phi):

you have to simplify your relation first: \[ 5(x + 2) \ge 7x + 2(x - 1)\] distribute the 5 on the left side (that means multiply 5 times each term inside the parens) and distribute the 2 on the right side. You get: \[ 5x+10 \ge 7x+2x-2 \] on the right side you have 7 x's add 2 more x's. that is 9x's \[ 5x+10 \ge 9x-2 \] now add -5x to both sides \[ 5x-5x+10 \ge 9x-5x-2 \]

OpenStudy (phi):

that simplifies to \[ 10 \ge 4x-2 \] add 2 to both sides to get \[ 12 \ge 4 x \] finally , divide both sides by 4 \[ 3 \ge x \] or \[ x \le 3 \] which means the same thing

OpenStudy (phi):

I guess if we just read the 4 choices, we could figure out that 3 of them mean the same thing. the one we *do not* want is the one that shows x is bigger than 3 (in other words we do not want to shade numbers bigger than 3)

OpenStudy (anonymous):

so than b would be the answer right? @phi

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@welshfella